### v1

# Heterogeneous model for superdiffusive movement of dense-core vesicles in C. elegans

Abstract
Transport of dense core vesicles (DCVs) in neurons is crucial for distributing molecules like neuropeptides and growth factors. We studied the experimental trajectories of dynein-driven directed movement of DCVs in the ALA neuron C. elegans over a duration of up to 6 seconds. We analysed the DCV movement in three strains of C. elegans: 1) with normal kinesin-1 function, 2) with reduced function in kinesin light chain 2 (KLC-2), and 3) a null mutation in kinesin light chain 1 (KLC-1). We find that DCVs move superdiffusively with displacement variance in all three strains with low reversal rates and frequent immobilization of DCVs. The distribution of DCV displacements fits a beta-binomial distribution with the mean and the variance following linear and quadratic growth patterns, respectively. We propose a simple heterogeneous random walk model to explain the observed superdiffusive retrograde transport behaviour of DCV movement. This model involves a random probability with the beta density for a DCV to resume its movement or remain in the same position.
keywords:
Heterogeneous anomalous diffusion, beta-binomial distribution, dense-core vesiclesIntroduction
The transport of dense-core vesicles (DCVs) is critical for maintaining cellular functions and organization, especially within neurons where efficient transport over long distances is necessary [1, 2]. This transport involves motor proteins kinesins and dyneins, which facilitate the movement of cargos along microtubules [3, 4]. Kinesins generally mediate anterograde transport, while dyneins drive retrograde transport [5, 6]. In neurons, DCVs transport diverse cargos, such as neuropeptides, monoamines, and neurotrophic factors [7, 8, 4]. DCVs play a crucial role in transporting various molecules essential for neuronal growth, signalling, learning, development, movement, and ageing [9, 7, 10, 11, 12]. Originating in the neuron soma, DCVs undergo fast axonal transport and release their content in response to specific stimuli, with fusion occurring at synaptic and non-synaptic sites [9, 3, 13, 14, 4].
We analyse DCV movement in the ALA neuron of nematode worms C. elegans. A notable characteristic of DCV motility in the ALA neuron is its bidirectional movement, which aligns with previous findings in various organisms [15, 13, 16]. The ALA neuron has a simpler structure compared to other multidendritic neurons and possesses one of the longest neurites in the nervous system of C. elegans [17]. Due to the narrow width of axons, moving DCVs occupy a significant portion of the axon diameter [18]. In wild-type neurons, stationary DCV accumulations are observed along the axon, with some DCVs moving, pausing, or stopping [3, 9, 13, 14, 4]. The ida-1 gene in C. elegans, an orthologue of mammalian type-1 diabetes auto-antigen proteins IA-2 and phogrin (IA-2), is crucial in neuroendocrine tissues for insulin secretion and glucose metabolism [19]. Transgenic ida-1::gfp nematodes were created by injecting N2 hermaphrodites with the ida-1p::ida-1::gfp DNA plasmid construct [19]. IDA-1 tagged with GFP is expressed in approximately 30 sensory neurons, including ALA, and is a component of neuropeptide-containing DCVs [19]. Confocal images of ida-1::gfp with GFP-tagged DCVs in the ALA neuron are shown in Figure 1. In the ALA neuron, DCVs exhibit predominantly unidirectional movement, with few reversals occurring at axon terminals or synapses [20].
Here, we focus on the dynein-driven DCV movement in three strains of C. elegans: 1) with normal kinesin-1 function (called ida-1::gfp), 2) with reduced function in kinesin light chain 2 (denoted as klc-2(rf) or KLC-2), and 3) a null mutation in kinesin light chain 1 (denoted as klc-1(-) or KLC-1). We use partial loss of function in KLC-2 because the null mutation is lethal. Even though the mutations are in kinesin-1, it has been shown that the inhibition of kinesin-1 commonly reduces the velocity for movement in both directions [15, 21], meaning that dynein movement is also different compared to non-mutated worms. We chose to study dynein movement as we have more data for retrograde movement than for anterograde. Additionally, KLC-1 and KLC-2 mutations has opposite effects on dynein movement, which makes the analysis more informative for our model. We observe that the variance of DCV displacements grows proportionally to . To describe this unusual superdiffusive behaviour, we develop a discrete random walk model with a stochastic probability of movement. Remarkably, this model explains the observed beta-binomial distribution of DCV displacements.
Traditional mathematical models of intracellular transport primarily involve Brownian motion concepts such as advection-diffusion, exclusion processes, and Brownian ratchet models [22, 23, 24]. However, the complexity of motor-driven motility often results in anomalous diffusion, which differs from Brownian motion [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. State-of-the-art machine learning methods are currently being developed for the analysis of anomalous diffusion in experimental datasets [39, 40, 41]. Interactions between kinesin and dynein motors, traditionally conceived as a “tug-of-war” between oppositely-directed motors attached to the same cargo [5], contribute to this complexity and are not yet fully understood. Intracellular transport is heterogeneous in both time and space, exhibiting varying non-Brownian behaviour [39, 36, 37]. Mathematical modelling of heterogeneous anomalous intracellular transport has been developed along non-Markovian random walk models [42], ensemble heterogeneity of diffusion coefficients [43], space-dependent diffusivity [44], and heterogeneity of jump probability [45]. A compartment model for DCVs transport in type II axonal terminals of Drosophila motoneurons on a large time scale has been developed [46].
In our recent paper [20], we studied the effect of kinesin-1 mutations on the transport of DCVs in ALA neurons in C. elegans. We showed how kinesin-1 mutations affect the lifespan and locomotion abilities of mutant worms. Here we focus on dynein-driven minus-end directed movement. The aim of this paper is to provide a simple data-driven model that offers insights into the heterogeneous movement of DCVs on a short time scale.
Results
Our goal was to determine the empirical distribution of displacements of dynein-driven DCVs and analyze its statistical characteristics such as the empirical mean and variance. Another aim was to develop a simple heterogeneous random walk model to describe these data. We find that the distribution of (Figure 2A) has an unimodal form with high dispersion, indicating heterogeneity of the populations of DCVs. While the empirical mean shows linear growth in time as expected, the variance unexpectedly shows a superdiffusive increase with time (Figure 2B, C). This behaviour drastically deviates from the expected linear increase of variance. This suggests that the movement of DCVs cannot be accurately described by a standard random walk model. As the variance of displacements of retrograde tracks resembles a quadratic behaviour, this indicates that the DCVs perform superdiffusive ballistic movement.
We observed that DCVs predominantly move unidirectionally with frequent immobilization of DCVs. To describe this random movement, we used a simple discrete random walk model. The DCV starts at zero, takes a small step in a small time interval on the right with the probability , or remains at the same position with the probability Then the probability mass function for the position of the DCV at the time is given by the binomial distribution To model the heterogeneity of the DCV population, we assume that the probability is a random parameter with the beta-density with , where is the beta function. Then one can obtain the well-known beta-binomial distribution[47, 48]
where and . The mean, and the variance, are
The essential feature of our heterogeneous DCV transport model is that the probability of successful movement is random. Assuming the randomness of this parameter, we use an indirect method to account for various factors influencing the movement of DCVs, including the size of DCVs, the level of activity of motor proteins, the structure of microtubules, etc. We assume that follows a beta distribution, with parameters and determining the shape of the distribution. The parameter quantifies the level of successful movement of DCVs. A higher value of means higher expected mobility of DCVs. The parameter measures the degree of failure in DCV movement. A higher value of corresponds to a situation where DCVs are immobilized, leading to a beta distribution skewed towards lower probabilities. By adjusting the values of and , one can generate a wide variety of beta distributions for effective modelling of the heterogeneity of DCV movement. This fits the general idea of the heterogeneity in biological particle movement, which refers to the variability observed in individual particle behaviour [37, 45]. It signifies that all DCVs exhibit different movement patterns or characteristics.
We fitted the beta-binomial distribution (1) to the empirical DCV dynein-driven displacement distribution by finding the optimal parameters and using Python. This was done through maximum likelihood estimation. The empirical displacement distribution for and and fitted beta-binomial distribution are shown in Figure 3. The resolution of the camera used for imaging is 1 pixel = 0.092 m, with a frame rate of 0.1563 seconds. Consequently, the time and the step size in (1) are chosen as s and m. The number of trials in the standard beta-binomial distribution , so at the time s in our random walk model DCV makes attempts to move on the right with the step size m. The number of tracks used to generate histograms (see Fig. 3) for 1.563 s and 3.126 s is as follows: ida-1::gfp- 851 and 549, klc-1(-); ida-1::gfp- 438 and 289, and klc-2(rf); ida-1::gfp- 1118 and 780 tracks respectively. Since the sample sizes used to generate histograms are relatively small, the heights of neighbouring bins exhibit high variability.
Figures 4 and 5 illustrate the excellent agreement between the theoretical mean and variance (equation (2)) and the empirical displacement data for each strain. It is quite surprising that a simple random walk model accurately captures the empirical ballistic superdiffusive behaviour of the displacement variance over a wide time span, from 1.5 to 5.5 seconds. This reflects the remarkable robustness of the heterogeneous mechanism leading to the beta-binomial distribution.
Discussion
In this study, we investigated the dynamics of the dynein-driven movement of a dense core vesicle (DCV) within the ALA neuron of three strains of C. elegans: those with normal kinesin-1 function, reduced function in kinesin light chain 2 (KLC-2), and a null mutation in kinesin light chain 1 (KLC-1). Our findings highlighted significant heterogeneity in DCV superdiffusive movement across all strains characterized by variances of displacements with quadratic growth in time () Our results demonstrated that in all three strains, the traditional random walk models were insufficient to describe the observed movement patterns of DCVs. The quadratic growth in variance suggests that DCV transport mechanisms involve a heterogeneous population of DCVs with different mobility properties. This is particularly evident in the differing impacts of the mutations on DCV motility. In strains with normal kinesin-1 function, DCV movement exhibited relatively stable and consistent transport patterns, consistent with efficient motor function. However, in the KLC-1 null mutation strain, we observed increased variability and a broader distribution of displacements. The KLC-2 reduced function strain showed the opposite effects, with slower DCV movement (Fig. 4). These observations highlight the differential roles of kinesin light chains in modulating motor protein interactions and cargo transport efficacy.
Numerous factors have an impact on the DCV’s mobility, including the presence and activity of motor proteins, the organization of microtubules, the size and nature of the cargo carried by DCVs, various proteins regulating vesicle trafficking, etc. To integrate these important factors into the transport model of DCV movement, one can use an indirect method by adjusting the parameters of the beta distribution. If we choose , the beta distribution is right-skewed, describing the situation of high mobility of DCVs. If the beta- distribution is left-skewed, which describes more immobilization of DCVs. For example, when it means complete immobilization. Small values of parameters and lead to a highly dispersed distribution describing high heterogeneity in DCV mobility. As an illustration, different sizes and natures of the cargo carried by DCVs can be taken into account in the beta-binomial distribution. Larger and heavily loaded DCVs move more slowly than smaller ones, therefore, the values of parameters for those vesicles should be , as we estimated for experimental trajectories for all mutants (see Figure 3).
The application of the beta-binomial distribution (1) provided an excellent fit for the DCV displacement data across all strains. This model accounts for the heterogeneity and discrete nature of the transport process, capturing the probabilistic essence of motor-driven movements better than conventional models. We identified the best-fit parameters and for each strain through optimisation, revealing strain-specific transport characteristics and underlying molecular mechanisms. The implications of our findings extend to understanding the molecular basis of intracellular transport heterogeneity. The superdiffusive quadratic growth in the variance of DCV displacements and the suitability of the beta-binomial distribution (1) underscore the complexity of motor protein dynamics and their regulation in a more realistic yet simple statistical model.
Methods
C. elegans strains
As a model organism, we use nematode worms Caenorhabditis elegans, (C. elegans). We used a strain expressing IDA-1::GFP, which allows for robust GFP (green fluorescent protein) labelling of DCVs in the ALA neuron [16] (Fig. 1). Kinesin-1 is a heterotetrameric molecular motor complex composed of two kinesin heavy chains (KHCs/KIF5) and two kinesin light chains (KLCs) [49]. To compare the motility of DCV in C. elegans with kinesin-1 mutations, we crossed the ida-1::gfp strain with mutants displaying reduced function of kinesin light chain 2 (klc-2(km11), referred to here as klc-2(rf)) [50], and a kinesin light chain 1 deletion mutant (klc-1(ok2609), referred to as klc-1(-)) [51]. In the genome of C. elegans, the unc-116 gene encodes a single kinesin heavy chain (KHC), and the klc-1 and klc-2 genes encode two kinesin light chains (KLCs) [50]. C. elegans strains were cultured and maintained according to standard protocols [52]. Worms were raised at 20°C on 6 cm plates (Greiner Bio-One) containing nematode growth media (NGM) seeded with OP-50 bacteria as food. The klc-1(ok2609) and klc-2(km11) strains were obtained from the Caenorhabditis Genetics Centre (CGC), while the ida-1::gfp strain was obtained from John Hutton’s laboratory at the University of Colorado Health Sciences Center, Denver, and kindly provided by Howard Davidson. Crosses of ida-1::gfp with klc-1(ok2609) and klc-2(km11) were performed as described [20].
Imaging
To observe DCV transport events in worms expressing IDA-1::GFP, we used time-lapse microscopy. Confocal imaging was conducted using a 3i inverted Spinning Disk Confocal Microscope equipped with a motorized stage for live-cell imaging. Images were captured with a CSU-X1 spinning disc confocal system (Yokogawa) on a Zeiss Axio-Observer Z1 microscope using various objectives, Prime 95B Scientific CMOS (1200 x1200 11 m pixels; backlit; 16-bit) camera (Photometrics) and motorised XYZ stage (ASI). Slidebook software (3i) was used for image acquisition.
Whole-worm images (Figure 1) were generated using Multiple XY Location Capture in SlideBook to produce montage images. Brightfield and fluorescence images were captured with the same camera, with LED brightfield illumination and a 100 ms exposure time. Fluorescence images were composed of multiple z-planes combined by maximum projection.
Videos of DCV movement were recorded using a 100x objective, with a single plane and a 156.3 ms frame interval for 500 frames. In each strain, videos were taken along the whole length of the ALA neuron as it was shown that distance from some does not affect the transport properties of DCVs [20]. Raw TIFF files were stabilized using the Image Stabilizer plugin for ImageJ, with adjustments made as needed for significant worm movement. Kymographs were generated using the ImageJ Multi Kymograph plugin, and tracks were identified using KymoButler [53]. The x and t positions for each track in every kymograph were obtained using the KymoButler code in Mathematica Version 14.0 (Wolfram Research, Inc.).
Analysis of the trajectories
We combined KymoButler results for each strain and performed analysis and visualization using custom Python code (available on GitHub upon publication). In instances where two x positions corresponded to the same t, indicating simultaneous particle occupancy, the average position was calculated. KymoButler outputs trajectories in pixels and timeframes format. We converted them to seconds and µm for further analysis (one timeframe corresponds to 0.1563 s and one pixel to 0.092 µm).
To isolate only dynein-driven retrograde tracks, we needed to exclude stationary and anterograde kinesin-driven tracks. This was achieved by selecting a threshold near the local minimum between retrograde and stationary peaks for each strain. Since the movement characteristics varied between strains, the thresholds were different. For the ida-1::gfp strain, we set the threshold at 1.6 the number of frames. For the klc-1(-);ida-1::gfp and klc-2(rf);ida-1::gfp strains, we chose a threshold of 1 the number of frames.
Acknowledgements
We extend our gratitude to Howard Davidson (University of Colorado Health Sciences Center, Denver) and the CGC for providing worm strains. The CGC is funded by the NIH Office of Research Infrastructure Programs (P40OD010440). A.G. would like to thank Gino Poulin (University of Manchester) for his assistance and guidance in working with C. elegans. The 3i spinning disk microscope used in this study was purchased through the University of Manchester Strategic Fund. Special thanks are due to Dr. Peter March for his help with confocal imaging. We are also grateful to Dmytro Chekunov for his assistance with Python and Mathematica coding. This work was funded by a Wellcome Trust PhD studentship awarded to AG (Grant No. 108867/Z/15/Z). SF, NK and VA were supported by EPSRC (Grant No. EP/V008641/1).
Author contributions statement
A.G. generated specific worm strains, imaging neurons with a spinning-disk confocal microscope, analysed DCV movement using KymoButler and custom Python code, segmented and calculated displacements, fitted a beta-binomial distribution to the data, prepared figures, and co-authoring the text of the paper. N.K. contributed to formal analysis, validation, investigation, and writing (review and editing). V.A. contributed to conceptualization, resources, data curation, supervision, funding acquisition, validation, methodology, project administration, and writing (review and editing). S.F. contributed to conceptualization, formal analysis, supervision, funding acquisition, validation, methodology, project administration, and writing (review and editing).
Additional information
Competing interests: No competing interests declared.